# Krystian Zawistowski 0009-0002-9589-9030 Samsung Research Poland Warsaw

### v1

@@ Abstract
LLM text decoding is key component for perceived LLM quality. We demonstrate two experiments showing that decoding methods could be improved by manipulation of token probabilities. First, we test few LLM on SummEval summary scoring dataset, to measure reading comprehension. We compare scores from greedy decoding to expected values over the next token distribution. We scale logits by large temperature to increase the entropy of scores. This allows strong improvement of performance on SummEval (in terms of correlations to human judgement). We see improvement from 6-8% to 13-28% for 7B Mistral and from 20%-46% to 37%-56% for Mixtral, beating GPT 4 0314 result on two metrics. Part of the gain seems related to positional bias. Secondly, we use probability-based tree sampling algorithm, to examine all most probable generations for given prompt.
@@ I Introduction
Generative LLM are trained on large text corpora as estimators of next token probability conditional on prior text. Then sampling from such probability distribution is performed, or token with largest probability is chosen (greedy decoding). Typically, one introduces parameter $T$ – called temperature. Let $l_{i}$ be the logit for $i$-th token. Then token probability is as follows:
@@
Question arises: what the T should be? Typically $T\in(0,1]$, with greedy decoding as limit in 0 and larger values corresponding to greater diversity (but also greater randomness). Research [1] shows that human generated text often does not correspond to modelled highest probability. Human choice of words is not guided by greatest probability, as “humans optimize against stating the obvious”. The author of this observation, H. P. Grice in [2] gave following example: suppose that I meet a man with a gas tank asking me to sell some gas and I answer “There’s gas station over the corner”. I said only a bit of information that doesn’t asnwer directly, while there’s lots of implicit information not being said (gas station is open, it seels gasoline and has it available and you can go there buy some). Such concepts might be pretty foreign to LLM: when we asked Mixtral Instruct “How to get gasoline in Fresno”, it gave us long instruction on finding gas station on a map, choosing best gas station, operating the pump safely, paying and so on, despite the fact that most of it is irrelevant to the problem of getting gasoline in Fresno specifically.
@@
Currently, greedy decoding is often used for label based QA [^1] For example in LM Evaluation Harness, standard set of LLM benchmarks [3]. We want to find whether relative probabilities of few most probable tokens are informative. We test whether greedy decoding can be outperformed by calculation of expected value. We evaluate our approach on SummEval [4] dataset. It contains 1600 article summaries with human annotations for relevance, fluency, coherence and consistency of a summary. We compare our result against known LLM based evaluations ([5], [6]). LLM is asked to evaluate relevance (or other feature) on Likert scale (from 1 to 5). We use MCQ prompts from [6], where LLM answer is A, B, C, D or E (A is 1 – worst, E is 5 – best). Let $p(A),p(B)...$ be probability of token “A”, “B”… Expected value score (that we use) is calculated as follows:
@@ II-B Results
Fluency 0,222 0,392 0,235 0,405
Relev. 0,457 0,555 0,464 0,564
Consist. 0,379 0,506 0,293 0,470
Coher. 0,428 0,485 0,388 0,438
Furthermore, strong improvements were produced for small and quantized LLM too. We evaluated 3 LLM, from 7B to 47B parameters and all metrics are consistently improved. We compare scores for quantized and float16 Mixtral Instruct in Table II. Surprisingly, quantized Mixtral performance is only slightly worse than float16 version.
@@
We see that summarization metrics, being relevant automated metrics for reading comprehension, strongly improve with the number of parameters. Also, quantized 4-bit LLM are very strong performers proportionally to their size and outperform similar size float16 models (for instance, quantized Mixtral, having 24GB in parameters and Mixture-of-Experts architecture strongly outperforms float16 SOLAR with 21 GB of parameters). For this reason quantized LLM might be viable, cost-effective option for RAG and other similar use cases. This phenomenon is similar to emergent abilities of LLM [11] where larger sizes lead to qualitative improvement in LLM performance. In this case too, quantized LLM [12] retain large portion of their emergent capabilities.
@@
Fluency 0,060 0,134 -0,045 0,061
Relev. 0,064 0,284 0,074 0,264
Consist. 0,061 0,252 0,076 0,249
Coher. 0,084 0,199 0,042 0,176
@@
Fluency 0,187 0,240 0,187 0,251
Relev. 0,192 0,427 0,165 0,364
Consist. 0,298 0,331 0,156 0,194
Coher. 0,305 0,362 0,200 0,267
@@
We performed this experiment for Mistral 7B for relevance evaluation and results can be found in Table V. Quite interestingly averaging out positional bias produces strong improvement for greedy decoding, while there’s no big difference for $E(s)$ decoding. Furthermore stronger improvement is found for Fp16 model, than for Int4 model. This suggests that gains from $E(s)$ method might be related to positional bias, but details of it are not clear without further research.
@@
These problems could be related to the use of softmax function in attention heads. Neural net limitations with respect to softmax and the rank of matrix under it are known and were brought to attention by [16] (which proposes high rank RNN). Similar problems might reemerge in case of transformers and attention, which use relatively small matrices for attention heads. For softmax it doesn’t matter whether signal is weak or strong, only whether it is the strongest among provided candidates. It is also true however, that weak signal supression can be learned by the attention head in the pretraining process, especially when bias matrices are added to $Q$ and $K$ – so it is impossible to tell more without further study.
@@ III Tree-based sampling
To further develop our hypothesis we propose an LLM inference analysis method that, for a given prompt, seeks to find all probable completions that could be generated by nuclueus sampling - to give complete, nearly deterministic picture, what LLM outputs could be for given prompt. As a foundation we use tree-search based sampling algorithm. We use priority queue mechanism, where most probable completions are evaluated first (like in Dijkstra algorithm). Tree sampling (a.k.a beam search) is broadly implemented approach [^2] Available in popular library Transformers [14]. in generative language models. Recently we saw very similar algorithm to ours [17] applied to compiler optimization (highest probability output produces superior compiler parametrization). Other controlled beam search techniques used for improved natural laguage generation can be found in [18], [19].
@@
  $\phi(t_{1}..t_{n})\rightarrow l_{i}$LLM model that maps token sequence to next-token log-probabilities}
@@
Outputs with their evaluated probabilities can be found in Table VI. We used nucleus sampling threshold $\hat{p}=0.9$ and temperature $T=2$ and we show outputs with $p>0.1\%$. Temperature is large, and reason is that for smaller temperatures only first, most probable output would be generated, while now it is generated with 73% probability. Output distribution is uneven, with top output 41 times more probable than second most probable output and very fat tailed with with about 23% of probability mass distributed among very unlikely outputs ($p<0.1\%$).
We didn not get any diversity of the paraphrase, despite asking for it explicitly, we got only two options: “I go by the name of Kennedy John” and “I go by the name of JFK”. At the same time LLM becomes unpredictable in the tail of the distribution as various additional comments follow after requested text. Output is thus not diverse and diversity we get provides little benefit: it might be a problem, when long unpredictable output follows the answer, evading usual stopping mechanisms of the inference (here we stop inference on two new lines in a row.)
@@
TABLE VI: Tree sampling algorithm output for Mixtral Instruct, completed senteces.
@@
Results point to causes of few problems of LLM that we believe to be fairly widespread in applications based on prompting LLM and parsing their output.
@@
    Repetitive output – LLM output might be identical most of time, as there is single completion with very large probability. LLM might sometimes provide little advantage compared to retrieval based or rule based system.
@@
These issues are rarely detected by commonly used accuracy-based benchmarks. There are some generic automated metrics such as MAUVE for diversity [20], but goal of our method is to analyze these issues in specific use cases and provide guidelines on configuration and further analysis. Our algorithm allows to analyze influence of modified prompts and system configuration on output probabilities, allowing greater degree of reliability and objectivity in the development, as different prompts, LLMs or sampling algorithm can be compared. On more general terms it seems that decoding heuristics like TopP seem to fail our expectationss, where entropy is extremely small or extremely large. Instead other approaches might be investigated: a) Scaling $T$ for entropy extremes of next-token probability distribution. b) Taboo sampling – tree sampling with penalty for token and substring repetition. c) Stopping generation for large entropy spikes (as that would likely result in unwanted output). This might allow new improvements in few important KPI for LLM, such as output diversity, controllability and safety. We used vLLM to implement our algorithm[13].
@@ IV Conclusions
We show reading comprehension metrics for 3 different LLM with sizes 7B, 10.7B and 47B with float16 (half precision) inference and 4 bit GPTQ quantization. Summarization metrics strongly improve with the number of parameters, and quantized 4-bit LLM are effective in proportion to their size (which is of interest for RAG on low-powered systems).

### v2

@@ Abstract
LLM text decoding is key component for perceived LLMs quality. We demonstrate two experiments showing that decoding methods could be improved by manipulation of token probabilities. First, we test few LLM on SummEval summary scoring dataset, to measure reading comprehension. We compare scores from greedy decoding to expected values over the next token distribution [^1] Source code released here: https://github.com/kzawisto/unused_information_llm. We scale logits by large temperature to increase the entropy of scores. This allows strong improvement of performance on SummEval (in terms of correlations to human judgement). We see improvement from 6-8% to 13-28% for 7B Mistral and from 20%-46% to 37%-56% for Mixtral, beating GPT 4 0314 result on two metrics. Part of the gain seems related to positional bias. Secondly, we use probability-based tree sampling algorithm, to examine all most probable generations for given prompt.
@@ I Introduction
Generative LLMs are trained on large text corpora as estimators of next token probability conditional on prior text. Then sampling from such probability distribution is performed, or token with largest probability is chosen (greedy decoding). Typically, one introduces parameter $T$ – called temperature. Let $l_{i}$ be the logit for $i$-th token. Then token probability is as follows:
@@
Question arises: what the T should be? Typically $T\in(0,1]$, with greedy decoding as limit in 0 and larger values corresponding to greater diversity (but also greater randomness). Research [1] shows that human generated text often does not correspond to modelled highest probability. Human choice of words is not guided by greatest probability, as “humans optimize against stating the obvious”. The author of this observation, H. P. Grice in [2] gave following example: suppose that I meet a man with a gas tank asking me to sell some gas and I answer “There’s gas station over the corner”. I said only a bit of information that doesn’t asnwer directly, while there’s lots of implicit information not being said (gas station is open, it seels gasoline and has it available and you can go there buy some). Such concepts might be pretty foreign to LLMs: when we asked Mixtral Instruct “How to get gasoline in Fresno”, it gave us long instruction on finding gas station on a map, choosing best gas station, operating the pump safely, paying and so on, despite the fact that most of it is irrelevant to the problem of getting gasoline in Fresno specifically.
@@
Currently, greedy decoding is often used for label based QA [^2] For example in LM Evaluation Harness, standard set of LLM benchmarks [3]. We want to find whether relative probabilities of few most probable tokens are informative. We test whether greedy decoding can be outperformed by calculation of expected value. We evaluate our approach on SummEval [4] dataset. It contains 1600 article summaries with human annotations for relevance, fluency, coherence and consistency of a summary. We compare our result against known LLM-based evaluations ([5], [6]). LLM is asked to evaluate relevance (or other feature) on Likert scale (from 1 to 5). We use MCQ prompts from [6], where LLM answer is A, B, C, D or E (A is 1 – worst, E is 5 – best). Let $p(A),p(B)...$ be probability of token “A”, “B”… Expected value score (that we use) is calculated as follows:
@@ II-B Results
Fluency 0.222 0.392 0.235 0.405
Relev. 0.457 0.555 0.464 0.564
Consist. 0.379 0.506 0.293 0.47
Coher. 0.428 0.485 0.388 0.438
Furthermore, strong improvements were produced for small and quantized LLM too. We evaluated 3 LLMs, from 7B to 47B parameters and all metrics are consistently improved. We compare scores for quantized and float16 Mixtral Instruct in Table II. Surprisingly, quantized Mixtral performance is only slightly worse than float16 version.
@@
We see that summarization metrics, being relevant automated metrics for reading comprehension, strongly improve with the number of parameters. Also, quantized 4-bit LLMs are very strong performers proportionally to their size and outperform similar size float16 models (for instance, quantized Mixtral, having 24GB in parameters and Mixture-of-Experts architecture strongly outperforms float16 SOLAR with 21 GB of parameters). For this reason quantized LLMs might be viable, cost-effective option for RAG and other similar use cases. This phenomenon is similar to emergent abilities of LLMs [11] where larger sizes lead to qualitative improvement in LLM performance. In this case too, quantized LLMs [12] retain large portion of their emergent capabilities.
@@
Fluency 0.06 0.134 -0.045 0.061
Relev. 0.064 0.284 0.074 0.264
Consist. 0.061 0.252 0.076 0.249
Coher. 0.084 0.199 0.042 0.176
@@
Fluency 0.187 0.24 0.187 0.251
Relev. 0.192 0.427 0.165 0.364
Consist. 0.298 0.331 0.156 0,194
Coher. 0.305 0.362 0.2 0.267
@@
We performed this experiment for Mistral 7B for relevance evaluation and results can be found in Table V. Quite interestingly averaging out positional bias produces strong improvement for greedy decoding, while there’s no big difference for $E(s)$ decoding. Furthermore stronger improvement is found for Fp16 model, than for Int4 model. Table VI shows few more experiments for Nous Hermes DPO Mistral 7B model, where prompt candidates are put either in ascending (A, B, C… order), reversed (E, D, C… order) or random order. NaN correlations indicate that the model had predicted identical result for all test examples. One new result here is that random order is remarkably bad, with many correlations dropping by 30% or more. From human point of view prompt says exact same thing, but this is not the same for LLMs, which cannot generalize when the structure is altered. Also, NousHermes Mistral, undergoing more extensive finetuning and alignment, outperforms Mistral Instruct on metrics related to logical reasoning, but underperforms on fluency and coherence.
This suggests that gains from $E(s)$ method might be related to positional bias, but details of it are not clear without further research.
@@
These problems could be related to the use of softmax function in attention heads. Neural net limitations with respect to softmax and the rank of matrix under it were brought to attention by [16] (which proposes high rank RNN). Similar problems might reemerge in case of transformers and attention, which use relatively small matrices for attention heads. For softmax it does not matter whether signal is weak or strong, only whether it is the strongest among provided candidates. It is also true however, that weak signal supression can be learned by the attention head in the pretraining process, especially when bias matrices are added to $Q$ and $K$ – so it is impossible to tell more without further study.
@@
TABLE VI: Pearson correlations to human judgement on SummEval for Mistral 7B Nous Hermes DPO fp16 - positional bias analysis.
Metric Greedy, ascending Greedy, reverse Greedy, random E(S), ascending E(S), reversed E(S), random
Relevance 0.32 0.36 0,092 0.44 0.42 0.2
Consistency 0.38 0.33 0.17 0.5 0.42 0.3
Fluency 0.097 NaN 0 0.12 0.0 0.015
Coherence 0.16 NaN 0.04 0.24 0.06 0.061
@@ III Tree-based sampling
To further develop our hypothesis we propose an LLM inference analysis method that, for a given prompt, seeks to find all probable completions that could be generated by nucleus sampling - to give complete, nearly deterministic picture, what LLM outputs could be for given prompt. As a foundation we use tree-search based sampling algorithm. We use priority queue mechanism, where most probable completions are evaluated first (like in Dijkstra algorithm). Tree sampling (a.k.a beam search) is broadly implemented approach [^3] Available in popular library Transformers [14]. in generative language models. Recently we saw very similar algorithm to ours [17] applied to compiler optimization (highest probability output produces superior compiler parametrization). Other controlled beam search techniques used for improved natural laguage generation can be found in [18], [19].
@@
  $\phi(t_{1}..t_{n})\rightarrow l_{i}$LLM that maps token sequence to next-token log-probabilities}
@@
Outputs with their evaluated probabilities can be found in Table VII. We used nucleus sampling threshold $\hat{p}=0.9$ and temperature $T=2$ and we show outputs with $p>0.1\%$. Temperature is large, and reason is that for smaller temperatures only first, most probable output would be generated, while now it is generated with 73% probability. Output distribution is uneven, with top output 41 times more probable than second most probable output and very fat tailed with with about 23% of probability mass distributed among very unlikely outputs ($p<0.1\%$).
We did not get any diversity of the paraphrase, despite asking for it explicitly, we got only two options: “I go by the name of Kennedy John” and “I go by the name of JFK”. At the same time LLM becomes unpredictable in the tail of the distribution as various additional comments follow after requested text. Output is thus not diverse and diversity we get provides little benefit: it might be a problem, when long unpredictable output follows the answer, evading usual stopping mechanisms of the inference (here we stop inference on two new lines in a row.)
@@
TABLE VII: Tree sampling algorithm output for Mixtral Instruct, completed senteces.
@@
Results point to causes of few problems of LLMs that we believe to be fairly widespread in applications based on prompting LLM and parsing their output.
@@
    Repetitive output – LLM output might be often identical, as there is single completion with very large probability. LLM might sometimes provide little advantage compared to retrieval based or rule based system.
@@
These issues are rarely detected by commonly used accuracy-based benchmarks. There are some generic automated metrics such as MAUVE for diversity [20], but goal of our method is to analyze these issues in specific use cases and provide guidelines on configuration and further analysis. Our algorithm allows to analyze influence of modified prompts and system configuration on output probabilities, allowing greater degree of reliability and objectivity in the development, as different prompts, LLMs or sampling algorithm can be compared. On more general terms it seems that decoding heuristics like TopP seem to fail our expectations, where entropy is extremely small or extremely large. Instead other approaches might be investigated: a) Scaling $T$ for entropy extremes of next-token probability distribution. b) Taboo sampling – tree sampling with penalty for token and substring repetition. c) Stopping generation for large entropy spikes (as that would likely result in unwanted output). This might allow new improvements in few important KPI for LLM, such as output diversity, controllability and safety.
@@ IV Conclusions
We show reading comprehension metrics for few different LLM with sizes 7B, 10.7B and 47B with float16 (half precision) inference and 4 bit GPTQ quantization. Summarization metrics strongly improve with the number of parameters, and quantized 4-bit LLM are effective in proportion to their size (which is of interest for RAG on low-powered systems).